# Least-Squares Neural Network (LSNN) Method for Linear Advection-Reaction Equation:

### v2

@@ Abstract
The least-squares ReLU neural network (LSNN) method was introduced and studied for solving linear advection-reaction equation with discontinuous solution in [4, 5]. The method is based on an equivalent least-squares formulation and employs ReLU neural network (NN) functions with $\lceil\log_{2}(d+1)\rceil+1$-layer representations for approximating solutions. In this paper, we show theoretically that the method is also capable of approximating non-constant jumps along discontinuous interfaces that are not necessarily straight lines. Numerical results for test problems with various non-constant jumps and interfaces show that the LSNN method with $\lceil\log_{2}(d+1)\rceil+1$ layers approximates solutions accurately with degrees of freedom less than that of mesh-based methods and without the common Gibbs phenomena along discontinuous interfaces.
@@ 1 Introduction
$\Gamma_{-}=\{{\bf x}\in\Gamma:\,\bm{\beta}({\bf x})\cdot\bm{n}({\bf x})<0\}$ (1)
with $\bm{n}({\bf x})$ being the unit outward normal vector to $\Gamma$ at ${\bf x}\in\Gamma$. Consider the linear advection-reaction equation
$\left\{\begin{array}[]{rccl}u_{\bm{\beta}}+\gamma\,u&=&f,&\text{ in }\,\Omega,\\[5.69054pt] u&=&g,&\text{ on }\,\Gamma_{-},\end{array}\right.$ (2)
@@
A major challenge in numerical simulation is that the solution of Eq. 2 is discontinuous along an interface because of a discontinuous inflow boundary condition, where the discontinuous interface can be the streamline from the inflow boundary. Traditional mesh-based numerical methods often exhibit oscillations near the discontinuity (called the Gibbs phenomena) and may not be extended to nonlinear hyperbolic conservation laws.
The least-squares ReLU neural network (LSNN) method for solving Eq. 2 with discontinuous solution was introduced and studied in [4, 5]. The method is based on an equivalent least-squares formulation studied in [2, 7] and employs ReLU neural network (NN) functions with $\lceil\log_{2}(d+1)\rceil+1$-layer representations for approximating the solution. The LSNN method is capable of automatically approximating the discontinuous solution since the free hyperplanes of ReLU NN functions adapt to the solution (see [3, 4, 5]). Compared to various adaptive mesh refinement (AMR) algorithms that locate the discontinuous interface through local mesh refinement (see, e.g., [6, 8, 11]), the LSNN method is much more effective in terms of the number of degrees of freedom.
Approximation properties of ReLU NN functions to step functions were recently studied in [4, 5]. In particular, we showed theoretically that two- or $\lceil\log_{2}(d+1)\rceil+1$-layer ReLU NN functions are necessary and sufficient to approximate a step function with any given accuracy $\varepsilon>0$ when the discontinuous interface is a hyperplane or general hyper-surface, respectively. This approximation property was used to establish a priori error estimates of the LSNN method.
The jump of the discontinuous solution of Eq. 2 is generally non-constant when the reaction coefficient $\gamma$ is non-zero. The main purpose of this paper is to establish a priori error estimates (see Theorem 3.3) for the LSNN method without making the assumption that the jump is constant. To this end, we decompose the solution as the sum of discontinuous and continuous parts (see Eq. 10), so that the discontinuous part of the solution can be described as a cylindrical surface on one subdomain and zero otherwise. Then we construct a continuous piecewise linear (CPWL) function with a sharp transition layer along the discontinuous interface to approximate the discontinuous part accurately. From [1, 5], we know that the CPWL function is a ReLU NN function $\mathbb{R}^{d}\to\mathbb{R}$ with a $\lceil\log_{2}(d+1)\rceil+1$-layer representation, from which it follows that the discontinuous part of the solution can be approximated by this class of functions for any prescribed accuracy. Then Theorem 3.3 follows.
The rest of the paper is organized as follows. In Section 2, we briefly review and discuss properties of ReLU NN functions and the LSNN method in [5]. Then theoretical convergence analysis is conducted in Section 3, showing that discretization error of the method for the problem mainly depends on the continuous part of the solution. Finally, to demonstrate the effectiveness of the method, we provide numerical results for test problems with various non-constant jumps in Section 4. Section 5 summarizes the work.
2 ReLU NN functions and the LSNN method
@@
This section briefly reviews properties of ReLU neural network (NN) functions and the least-squares ReLU neural network (LSNN) method in [5]. A function $\mathcal{N}:\mathbb{R}^{d}\to\mathbb{R}^{c}$ is called a ReLU neural network (NN) function if it can be expressed as a composition of functions
$N^{(L)}\circ\cdots\circ N^{(2)}\circ N^{(1)}\text{ with }L>1,$ (3)
@@
where $N^{(l)}:\mathbb{R}^{n_{l-1}}\to\mathbb{R}^{n_{l}}$ ($n_{0}=d$, $n_{L}=c$) is affine linear when $l=L$, and affine linear with the rectified linear unit (ReLU) activation function $\sigma$ applied to each component when $11\leq l\leq L-$ 1. Each affine linear function takes the form $\bm{\omega}^{(l)}\mathbf{x}-\mathbf{b}^{(l)}$ for $\mathbf{x}\in\mathbb{R}^{n_{l-1}}$ where $\bm{\omega}^{(l)}\in\mathbb{R}^{n_{l}\times n_{l-1}}$, $\mathbf{b}^{(l)}\in\mathbb{R}^{n_{l}}$ are weight and bias matrices, respectively. For a given positive integer $n$, denote the collection of all ReLU NN functions from $\mathbb{R}^{d}$ to $\mathbb{R}$ that have representations with depth $L$ and total number of hidden neurons $n(=n_{1}+\cdots+n_{L-1})$ by $\mathcal{M}(d,1,L,n)$ (1 being the output dimension), and the collection of all ReLU NN functions from $\mathbb{R}^{d}$ to $\mathbb{R}$ with $L$-layer representations by $\mathcal{M}(d,1,L)$. Then
$\mathcal{M}(d,1,L)=\bigcup_{n\in\mathbb{N}}\mathcal{M}(d,1,L,n).$ (4)
@@
The collection of all continuous piecewise linear (CPWL) functions on $\mathbb{R}^{d}$ is equal to ${\cal M}(d,1,\lceil\log_{2}(d+1)\rceil+1)$, i.e., the collecion of all ReLU NN functions from $\mathbb{R}^{d}$ to $\mathbb{R}$ that have representations with depth $\lceil\log_{2}(d+1)\rceil+1$.
@@
Finally, breaking hyperplanes are depicted in Figs. 2, 3, 4, 5, 6, and 7 to better understand the graphs of ReLU NN function approximations using domain partitions (on each element in a given partition, the ReLU NN function is affine linear; see [5]). More specifically, the $l^{\text{th}}$- (hidden) layer breaking hyperplanes of a given ReLU NN function (with output dimension 1) representation as in (3) are defined as the union of the zero sets of the functions
  • $\bm{\omega}^{(1)}\mathbf{x}-\mathbf{b}^{(1)}$ when $l=1$ ($n_{1}$ functions),
  • $\bm{\omega}^{(l)}(N^{(l-1)}\circ\cdots\circ N^{(2)}\circ N^{(1)}(\mathbf{x}))-\mathbf{b}^{(l)}$ when $2\leq l<L$ ($n_{l}$ functions).
@@
$\mathcal{L}(v;{\bf f})=\|v_{\bm{\beta}}+\gamma\,v-f\|_{0,\Omega}^{2}+\|v-g\|_{-\bm{\beta}}^{2},$ (5)
@@
$\mathcal{L}(u;{\bf f})=\min_{\small v\in V_{\bm{\beta}}}\mathcal{L}(v;{\bf f}),$ (6)
@@
Then the corresponding LS and discrete LS approximations are, respectively, to find
$u_{{}_{N}}\in{\cal M}(d,1,L,n)$ such that
$\mathcal{L}\big{(}u_{{}_{N}};{\bf f}\big{)}=\min\limits_{v\in{\cal M}(d,1,L,n)}\mathcal{L}\big{(}v;{\bf f}\big{)},$ (7)
@@
$\mathcal{L}_{{}_{\small{\cal T}}}\big{(}{u}^{{}_{N}}_{{}_{\small{\cal T}}};{\bf f}\big{)}=\min\limits_{v\in{\cal M}(d,1,L,n)}\mathcal{L}_{{}_{\small{\cal T}}}\big{(}v;{\bf f}\big{)},$ (8)
where $\mathcal{L}_{{}_{\small{\cal T}}}\big{(}v;{\bf f}\big{)}$ is the discrete LS functional (see [4]).
@@ 3 Error estimates
$\chi({\bf x})=\left\{\begin{array}[]{rl}a({\bf x}),&{\bf x}\in\Omega_{1i},\\[5.69054pt] 0,&{\bf x}\in\Omega_{2i};\end{array}\right.$ (9)
@@
${\left|\kern-1.07639pt\left|\kern-1.07639pt\left|\chi-p_{i}\right|\kern-1.07639pt\right|\kern-1.07639pt\right|}_{\bm{\beta}}\leq D_{1}\sqrt{\varepsilon_{2}}+D_{2}\sqrt{\varepsilon_{3}}.$ (11)
@@
${\left|\kern-1.07639pt\left|\kern-1.07639pt\left|u-u_{{}_{N}}\right|\kern-1.07639pt\right|\kern-1.07639pt\right|}_{\bm{\beta}}\leq C\,\left(\sqrt{\varepsilon_{2}}+\sqrt{\varepsilon_{3}}+\inf_{v\in{\cal M}(d,n-\hat{n})}{\left|\kern-1.07639pt\left|\kern-1.07639pt\left|\hat{u}-v\right|\kern-1.07639pt\right|\kern-1.07639pt\right|}_{\bm{\beta}}\right),$ (12)
@@
$\begin{split}{\left|\kern-1.07639pt\left|\kern-1.07639pt\left|u-u^{{}_{N}}_{{}_{\cal T}}\right|\kern-1.07639pt\right|\kern-1.07639pt\right|}_{\bm{\beta}}\leq\ &C_{1}\,\left(\big{|}({\cal L}-{\cal L}_{{}_{\cal T}})(u_{{}_{N}}-u_{{}_{\cal T}}^{{}_{N}},{\bf 0})\big{|}+\big{|}({\cal L}-{\cal L}_{{}_{\cal T}})(u-u_{{}_{N}},{\bf 0})\big{|}\right)^{1/2}\\ &+C_{2}\,\left(\sqrt{\varepsilon_{2}}+\sqrt{\varepsilon_{3}}+\inf_{v\in{\cal M}(d,n-\hat{n})}{\left|\kern-1.07639pt\left|\kern-1.07639pt\left|\hat{u}-v\right|\kern-1.07639pt\right|\kern-1.07639pt\right|}_{\bm{\beta}}\right).\end{split}$ (13)
@@ 4 Numerical experiments
In this section, we present numerical results for both two- and three-dimensional test problems with constant, piecewise constant, or variable advection velocity fields. The discrete LS functional was minimized by the Adam optimization algorithm [9] on a uniform mesh with mesh size $h=10^{-2}$. The directional derivative $v_{\bm{\beta}}$ was approximated by the backward finite difference quotient multiplied by $|{\bm{\beta}}|$
$v_{\bm{\beta}}({\bf x}_{{}_{K}})\approx|{\bm{\beta}}|\frac{v({\bf x}_{{}_{K}})-v\big{(}{\bf x}_{{}_{K}}-\rho\bar{\bm{\beta}}({\bf x}_{{}_{K}})\big{)}}{\rho},$ (14)
@@
All of the test problems are defined on the domain $\Omega=(0,1)^{2}$ or $(0,1)^{3}$ with $\gamma=1$ ($f=1$ for the first three and the last test problems, and $f=0$ for the remaining test problems).
@@ 4.1 A problem with a constant advection velocity field
The advective velocity field is a constant field given by
$\bm{\beta}(x,y)=(0,1),\,\,(x,y)\in\Omega.$ (15)
The inflow boundary and the inflow boundary condition are given by
$\Gamma_{-}$ $=$ $\{(x,0):x\in(0,1)\}$
and $\mbox{and }\,\,g(x,y)$ $=$ $\left\{\begin{array}[]{rl}1,&(x,y)\in\Gamma^{1}_{-}\equiv\{(x,0):x\in(0,1/2)\},\\[5.69054pt] 2,&(x,y)\in\Gamma^{2}_{-}=\Gamma_{-}\setminus\Gamma_{-}^{1},\end{array}\right.$
respectively. The exact solution of this test problem is
$u(x,y)=\left\{\begin{array}[]{rl}1,&(x,y)\in\Omega_{1}=\{(x,y)\in\Omega:x<1/2\},\\[5.69054pt] 1+e^{-y},&(x,y)\in\Omega\setminus\Omega_{1}.\end{array}\right.$ (17)
50000 iterations were implemented with 2–20–20–1 ReLU NN functions. The numerical results are presented in Figs. 2 and 1. The traces (Fig. 2(b)) of the exact and numerical solutions on the plane $y=0.5$ show no difference or oscillation. The exact solution (Fig. 2(c)), which has a non-constant jump along the vertical interface (Fig. 2(a)) is accurately approximated by a 3-layer ReLU NN function (Figs. 2(d) and 1). We note that the solution of this test problem takes the same form as $\chi(\mathbf{x})$ in (9), which was approximated by a CPWL function constructed by partitioning the domain into rectangles stacking on top of each other. It appears from Fig. 2(e) that the 3-layer ReLU NN function approximation has a similar partition, and the second-layer breaking hyperplanes were generated for approximating the jump along the discontinuous interface and the non-constat part of the solution, which is consistent with our theoretical analysis on the convergence of the method.

### v3

@@ Abstract
The least-squares ReLU neural network (LSNN) method was introduced and studied for solving linear advection-reaction equation with discontinuous solution in [4, 5]. The method is based on an equivalent least-squares formulation and [5] employs ReLU neural network (NN) functions with $\lceil\log_{2}(d+1)\rceil+1$-layer representations for approximating solutions. In this paper, we show theoretically that the method is also capable of accurately approximating non-constant jumps along discontinuous interfaces that are not necessarily straight lines. Theoretical results are confirmed through multiple numerical examples with $d=2,3$ and various non-constant jumps and interface shapes, showing that the LSNN method with $\lceil\log_{2}(d+1)\rceil+1$ layers approximates solutions accurately with degrees of freedom less than that of mesh-based methods and without the common Gibbs phenomena along discontinuous interfaces having non-constant jumps.
@@ 1 Introduction
For decades, extensive research has been conducted on numerical methods for linear advection-reaction equations to develop precise and efficient numerical schemes. A major challenge in numerical simulation is that the solution of the equation is discontinuous along an interface because of a discontinuous inflow boundary condition, where the discontinuous interface can be the streamline from the inflow boundary. Traditional mesh-based numerical methods often exhibit oscillations near the discontinuity (called the Gibbs phenomena), which are not suitable for many applications and may not be extended to nonlinear hyperbolic conservation laws.
Recently, the application of neural networks (NNs) for solving partial differential equations have achieved significant accomplishments. For the linear advection-reaction problem, the least-squares ReLU neural network (LSNN) method was introduced and studied in [4, 5]. The method is based on an equivalent least-squares formulation studied in [2, 6] and [5] employs ReLU neural network (NN) functions with $\lceil\log_{2}(d+1)\rceil+1$-layer representations for approximating the solution. The LSNN method is capable of automatically approximating the discontinuous solution accurately since the free hyperplanes of ReLU NN functions adapt to the solution (see [3, 4, 5]). Moreover, for problems with unknown locations of discontinuity interfaces, it is quite easy to see that the LSNN method uses much fewer number of degrees of freedom than the mesh-based methods (see, e.g., [3, 4]).
Approximation properties of ReLU NN functions to step functions were examined and employed in [4, 5]. In particular, it was shown theoretically that two- or $\lceil\log_{2}(d+1)\rceil+1$-layer ReLU NN functions are necessary and sufficient to approximate a step function with any given accuracy $\varepsilon>0$ when the discontinuous interface is a hyperplane or general hyper-surface, respectively. This approximation property was used to establish a priori error estimates of the LSNN method for the linear advection-reaction problem.
The jump of the discontinuous solution of the problem, however, is generally non-constant when the reaction coefficient is non-zero. The main purpose of this paper is to establish a priori error estimates (see Theorem 3.3) for the LSNN method without making the assumption that the jump is constant as in [5]. To this end, we decompose the solution as the sum of discontinuous and continuous parts (see Eq. 10), so that the discontinuous part of the solution can be described as a cylindrical surface on one subdomain and zero otherwise. Then we construct a continuous piecewise linear (CPWL) function with a sharp transition layer along the discontinuous interface to approximate the discontinuous piecewise cylindrical surface accurately. From [10, 11, 1, 5], we know that the CPWL function is a ReLU NN function $\mathbb{R}^{d}\to\mathbb{R}$ with a $\lceil\log_{2}(d+1)\rceil+1$-layer representation, from which it follows that the discontinuous part of the solution can be approximated by this class of functions for any prescribed accuracy. Then Theorem 3.3 follows.
The rest of the paper is organized as follows. In Section 2, we introduce the linear advection-reaction problem, and briefly review and discuss properties of ReLU NN functions and the LSNN method in [5]. Then theoretical convergence analysis is conducted in Section 3, showing that discretization error of the method for the problem mainly depends on the continuous part of the solution. Finally, to demonstrate the effectiveness of the method, we provide numerical results for test problems with various non-constant jumps in Section 4. Section 5 summarizes the work.
2 Problem formulation and the LSNN method
@@
(1) $\Gamma_{-}=\{{\bf x}\in\Gamma:\,\bm{\beta}({\bf x})\cdot\bm{n}({\bf x})<0\}$
where $\bm{n}({\bf x})$ is the unit outward normal vector to $\Gamma$ at ${\bf x}\in\Gamma$. Consider the linear advection-reaction equation
(2) $\left\{\begin{array}[]{rccl}u_{\bm{\beta}}+\gamma\,u&=&f,&\text{ in }\,\Omega,\\[5.69054pt] u&=&g,&\text{ on }\,\Gamma_{-},\end{array}\right.$
@@
For the convenience of the reader, this section briefly reviews properties of ReLU neural network (NN) functions and the least-squares ReLU neural network (LSNN) method in [5]. A function $\mathcal{N}:\mathbb{R}^{d}\to\mathbb{R}^{c}$ is called a ReLU neural network (NN) function if it can be expressed as a composition of functions
(3) $N^{(L)}\circ\cdots\circ N^{(2)}\circ N^{(1)}\text{ with }L>1,$
@@
where $N^{(l)}:\mathbb{R}^{n_{l-1}}\to\mathbb{R}^{n_{l}}$ ($n_{0}=d$, $n_{L}=c$) is affine linear when $l=L$, and affine linear with the rectified linear unit (ReLU) activation function $\sigma$ applied to each component when $11\leq l\leq L-$ 1. Each affine linear function takes the form $\bm{\omega}^{(l)}\mathbf{x}-\mathbf{b}^{(l)}$ for $\mathbf{x}\in\mathbb{R}^{n_{l-1}}$ where $\bm{\omega}^{(l)}\in\mathbb{R}^{n_{l}\times n_{l-1}}$, $\mathbf{b}^{(l)}\in\mathbb{R}^{n_{l}}$ are weight and bias matrices, respectively. For $n\in\mathbb{N}$, denote the collection of all ReLU NN functions from $\mathbb{R}^{d}$ to $\mathbb{R}$ with depth $L$ and the number of hidden neurons $n(=n_{1}+\cdots+n_{L-1})$ by $\mathcal{M}(d,1,L,n)$ (1 being the output dimension), and the collection of all ReLU NN functions from $\mathbb{R}^{d}$ to $\mathbb{R}$ with depth $L$ by $\mathcal{M}(d,1,L)$. Then we have
(4) $\mathcal{M}(d,1,L)=\bigcup_{n\in\mathbb{N}}\mathcal{M}(d,1,L,n).$
@@
The collection of all continuous piecewise linear (CPWL) functions on $\mathbb{R}^{d}$ is equal to ${\cal M}(d,1,\lceil\log_{2}(d+1)\rceil+1)$, i.e., the collection of all ReLU NN functions from $\mathbb{R}^{d}$ to $\mathbb{R}$ that have representations with depth $\lceil\log_{2}(d+1)\rceil+1$.
@@
As in [5], breaking hyperplanes are depicted in Figs. 2, 3, 4, 5, 6, and 7 to better understand the graphs of ReLU NN function approximations using domain partitions (on each element in a given partition, the ReLU NN function is affine linear; see [5]). More specifically, the $l^{\text{th}}$- (hidden) layer breaking hyperplanes of a given ReLU NN function (with output dimension 1) representation as in (3) are defined as the zero sets of the layer without the activation function: when $l=1$, $\mathbf{w}_{i}^{(1)}\mathbf{x}-b_{i}^{(1)}=0$, and when $2\leq l<L$,
$\mathbf{w}_{i}^{(l)}\left(N^{(l-1)}\circ\cdots\circ N^{(2)}\circ N^{(1)}(\mathbf{x})\right)-b_{i}^{(l)}=0,$
where
$\bm{\omega}^{(l)}=(\mathbf{w}_{1}^{(l)},\ldots,\mathbf{w}_{n_{l}}^{(l)})^{T}\in\mathbb{R}^{n_{l}\times n_{l-1}},\quad\text{and}\quad\mathbf{b}^{(l)}=(b_{1}^{(l)},\ldots,b_{n_{l}}^{(l)})^{T}.$
@@
(5) $\mathcal{L}(v;{\bf f})=\|v_{\bm{\beta}}+\gamma\,v-f\|_{0,\Omega}^{2}+\|v-g\|_{-\bm{\beta}}^{2},$
@@
(6) $\mathcal{L}(u;{\bf f})=\min_{\small v\in V_{\bm{\beta}}}\mathcal{L}(v;{\bf f}),$
@@
Then the corresponding LS and discrete LS approximations are, respectively, to find $u_{{}_{N}}\in{\cal M}(d,1,L,n)$ such that
(7) $\mathcal{L}\big{(}u_{{}_{N}};{\bf f}\big{)}=\min\limits_{v\in{\cal M}(d,1,L,n)}\mathcal{L}\big{(}v;{\bf f}\big{)},$
@@
(8) $\mathcal{L}_{{}_{\small{\cal T}}}\big{(}{u}^{{}_{N}}_{{}_{\small{\cal T}}};{\bf f}\big{)}=\min\limits_{v\in{\cal M}(d,1,L,n)}\mathcal{L}_{{}_{\small{\cal T}}}\big{(}v;{\bf f}\big{)},$
where $\mathcal{L}_{{}_{\small{\cal T}}}\big{(}v;{\bf f}\big{)}$ is the discrete LS functional (see [4, 5]).
@@ 3 Error estimates
(9) $\chi({\bf x})=\left\{\begin{array}[]{rl}a({\bf x}),&{\bf x}\in\Omega_{1i},\\[5.69054pt] 0,&{\bf x}\in\Omega_{2i};\end{array}\right.$
@@
(11) ${\left|\kern-1.07639pt\left|\kern-1.07639pt\left|\chi-p_{i}\right|\kern-1.07639pt\right|\kern-1.07639pt\right|}_{\bm{\beta}}\leq D_{1}\sqrt{\varepsilon_{2}}+D_{2}\sqrt{\varepsilon_{3}}.$
@@
(12) ${\left|\kern-1.07639pt\left|\kern-1.07639pt\left|u-u_{{}_{N}}\right|\kern-1.07639pt\right|\kern-1.07639pt\right|}_{\bm{\beta}}\leq C\,\left(\sqrt{\varepsilon_{2}}+\sqrt{\varepsilon_{3}}+\inf_{v\in{\cal M}(d,n-\hat{n})}{\left|\kern-1.07639pt\left|\kern-1.07639pt\left|\hat{u}-v\right|\kern-1.07639pt\right|\kern-1.07639pt\right|}_{\bm{\beta}}\right),$
@@
(13) $\begin{split}{\left|\kern-1.07639pt\left|\kern-1.07639pt\left|u-u^{{}_{N}}_{{}_{\cal T}}\right|\kern-1.07639pt\right|\kern-1.07639pt\right|}_{\bm{\beta}}\leq\ &C_{1}\,\left(\big{|}({\cal L}-{\cal L}_{{}_{\cal T}})(u_{{}_{N}}-u_{{}_{\cal T}}^{{}_{N}},{\bf 0})\big{|}+\big{|}({\cal L}-{\cal L}_{{}_{\cal T}})(u-u_{{}_{N}},{\bf 0})\big{|}\right)^{1/2}\\ &+C_{2}\,\left(\sqrt{\varepsilon_{2}}+\sqrt{\varepsilon_{3}}+\inf_{v\in{\cal M}(d,n-\hat{n})}{\left|\kern-1.07639pt\left|\kern-1.07639pt\left|\hat{u}-v\right|\kern-1.07639pt\right|\kern-1.07639pt\right|}_{\bm{\beta}}\right).\end{split}$
@@ 4 Numerical experiments
In this section, we demonstrate the performance of the LSNN method across different settings, incorporating various non-constant jumps. The discrete LS functional was minimized by the Adam optimization algorithm [7] on a uniform mesh with mesh size $h=10^{-2}$. As in [5], the directional derivative $v_{\bm{\beta}}$ was approximated by the backward finite difference quotient multiplied by $|{\bm{\beta}}|$
(14) $v_{\bm{\beta}}({\bf x}_{{}_{K}})\approx|{\bm{\beta}}|\frac{v({\bf x}_{{}_{K}})-v\big{(}{\bf x}_{{}_{K}}-\rho\bar{\bm{\beta}}({\bf x}_{{}_{K}})\big{)}}{\rho},$
@@
All the test problems are defined on the domain $\Omega=(0,1)^{2}$ or $(0,1)^{3}$ with $\gamma=1$ ($f=1$ for the first three and the last test problems, and $f=0$ for the remaining test problems).
@@ 4.1 A problem with a constant advection velocity field
The first test problem has the constant advective velocity field $\bm{\beta}(x,y)=(0,1),\,\,(x,y)\in\Omega$, and the inflow boundary of the problem is $\Gamma_{-}=\{(x,0):x\in(0,1)\}$. The inflow boundary condition is given by
$g(x,y)=\left\{\begin{array}[]{rl}1,&(x,y)\in\Gamma^{1}_{-}\equiv\{(x,0):x\in(0,1/2)\},\\[5.69054pt] 2,&(x,y)\in\Gamma^{2}_{-}=\Gamma_{-}\setminus\Gamma_{-}^{1}.\end{array}\right.$
The exact solution of this test problem is
$u(x,y)=\left\{\begin{array}[]{rl}1,&(x,y)\in\Omega_{1}=\{(x,y)\in\Omega:x<1/2\},\\[5.69054pt] 1+e^{-y},&(x,y)\in\Omega\setminus\Omega_{1}.\end{array}\right.$
The LSNN method was implemented with 50000 iterations for 2–20–20–1 ReLU NN functions. We report the numerical results in Figs. 2 and 1. The traces (Fig. 2(b)) of the exact and numerical solutions on the plane $y=0.5$ show no difference or oscillation. The exact solution (Fig. 2(c)), which has a non-constant jump along the vertical interface (Fig. 2(a)) is accurately approximated by a 3-layer ReLU NN function (Figs. 2(d) and 1). We note that the solution of this test problem takes the same form as $\chi(\mathbf{x})$ in (9), which was approximated by a CPWL function constructed by partitioning the domain into rectangles stacking on top of each other. It appears from Fig. 2(e) that the 3-layer ReLU NN function approximation has a similar partition, and the second-layer breaking hyperplanes were generated for approximating the jump along the discontinuous interface and the non-constant part of the solution, which is consistent with our theoretical analysis on the convergence of the method.